# Scalable Offline Metrics for Autonomous Driving

### v1

@@ IV Experiments
We first describe our experimental setup and model training and testing procedure. We then leverage the setup to perform comprehensive experiments in simulation over various towns, ambient settings, and traffic infractions using the CARLA simulation [34]. However, while most online performance studies for autonomous driving are performed in simulation, this may bias results, i.e., due to appearance and physics artifacts. To ensure our findings generalize to real-world models, we thus also incorporate analysis using a small-scale platform in the real-world.
@@ IV-A Experimental Setup
Simulation Benchmark: We follow the standard Longest6 testing setup  [41, 9]. Longest6 benchmark comprises 36 routes with an average route length of 1.5km, which goes beyond the prior study on CARLA’s Town02 by Codevilla [28]. We leverage a privileged rule-based expert to generate a large training and validation dataset, where trained policies are evaluated over unseen weather and routes in training.
@@
Real-World Benchmark: Online evaluation of driving policies is generally done in simulation [41, 67, 10]. To ensure our findings generalize to the real-world, we utilize a safe, small-scale platform. This platform is based on a commercial RC car (Traxxas X-Maxx [68]). We set up a Jetson Orin and a wide-angle camera to capture observations and control the steering and speed of the car using custom scripts and a joystick controller. We note that we do not leverage LiDAR in our real-world platform. We collect a training set of $32,356$ frames on various sidewalks, times of day, and weather. We evaluate on an uneven road that has three right turns, two left turns, and three pedestrians crossing (example routes are shown in Fig. 2). The platform can travel at speeds of up to 5050MPH. Online evaluations are safely done with an emergency software stop. In addition to the long route evaluation, we also perform targeted evaluations with clearly defined scenarios over traffic lights, vehicles, and object crossings.
@@
Model Training Strategies: In the simulation, our baseline model is TransFuser [9], which provides a state-of-the-art conditional imitation learning model. The model is a multi-modal policy with two inputs: RGB image and LiDAR. The features of these modalities are fused together by a transformer-based architecture with a cross-modal attention mechanism to predict output actions. To ensure our findings are applicable to a broad range of models, we train a large set of model variations by modifying the model input (e.g., RGB only vs. RGB and LiDAR) and backbone (two ConvNext [69], small and tiny, and nine RegNet [70], from RegNetY-200MF to RegNetY-8.0GF). We train each model (input and backbone combination) for 40 epochs while also saving the checkpoints every 10 epochs to sample additional model variations. Applying test-time dropouts provides another means of sampling models and covering the space of all possible policies. Moreover, we sample models using test-time dropout for values between 10% to 30% with increments of 10%. The Pearson correlation coefficient is then computed over all sampled models. Results are presented across 82 models, each varying in backbones, epochs, inputs, and test-time dropout rates. For real-world models, in the interest of optimizing for real-time performance, we target smaller and simpler architectures motivated by prior works from Codevilla et al. [12], Hawke et al. [71] and Bojarski et al. [38] with a baseline of a five-layer CNN as well as variants of RegNetY as feature extractor backbones. In the real-world experiment, we train an end-to-end policy model solely using RGB input, as it provides an easy and scalable implementation for researchers to reproduce and build upon in the future (i.e., without LiDAR).
@@
Figure 3: Driving Score Correlation Analysis in Simulation. The plot shows updated correlation for offline metrics, including TRE, QCE and PDM, the most successful reported metrics by prior research [28, 61], as well as widely employed metrics such as waypoint FDE (Final waypoint Displacement Error). Given our updated evaluation setup with complex traffic scenarios and diverse models, correlations are low overall, besides the introduced uncertainty-weighted (UW) version. Each disc shows one model sampled from a certain epoch, backbone, input (with or without LiDAR), and a test-time dropout rate used to obtain coverage of models with varying performances (see further explanation in Sec. IV-A). The radius of each marker is proportional to the percent of test-time dropout used (this sampling is independent of our proposed metric).
@@ IV-B Results in Simulation
Given the more extensive benchmark and model sampling strategies, i.e., with dropout over model weights for broader coverage of the performance spectrum, we revisit the analysis of Codevilla et al. [28] (only analyzed steering-based metrics, while fixing other action values based on the ground-truth). We also leverage the Driving Score, as it provides a holistic measure for driving quality, i.e., beyond Success Rate, as well as finer-grained online metrics.
@@
Revisiting Correlation in Simulation: As shown in Fig. 3, we show an overall poor correlation (both for Driving Score and Success Rate, isolated in Fig. 4) for standard offline metrics. Specifically, QCE and TRE, which are designed for higher correlation with the online Success Rate metric, show poor correlation (0.22 and 0.56, respectively) to more holistic online performance metrics such as the Driving Score. As shown in Fig. 3, QCE and TRE even perform worse than FDE (waypoint-based final displacement error), which is a standard offline metric for high-performing systems today. We hypothesize that the reason these were successful is due to the highly simplistic test settings in Codevilla et al. [28] (Town02, not considering collisions or infractions). However, we observe that our proposed UWE achieves the highest correlation within this challenging benchmark, $0.69$, $0.08$ higher than prior offline metrics. We only show selected metrics, as the remaining analyzed offline metrics performed with worse correlation. We study nine online metrics to get a complete understanding of the strengths and weaknesses of these metrics and identify key aspects that help choose the most promising metric. Our UWE demonstrates a stronger correlation with all online metrics and significantly outperforms other common offline metric baselines, as shown in Fig. 4. We also compare our UWE metric with the PDM (Predictive Driver Model) score from NAVSIM [61], as shown in Fig. 3, using the publicly available code. As shown in Fig. 3, the PDM score is lower in our more realistic CARLA simulation compared to its performance on nuPlan. Specifically, in settings with more complex physics and reactivity, UWE’s correlation exhibits significantly higher fidelity than the PDM score. We note that this is likely because the PDM score makes assumptions about agent reactivity and environmental information (i.e., being overly accommodating). In contrast, UWE does not require such assumptions. As such, our metric provides a simple yet effective and scalable alternative potentially applicable to a broader range of situations with various agent dynamics.
@@ IV-D UWE Ablation
We further investigate the reliability of the uncertainty estimates obtained via Monte Carlo Dropout. When we replace the dropout mechanism with a model ensemble and observe an increase in correlation from 0.69 to 0.80 in the Pearson coefficient. While this aligns with prior literature [37], dropout-based inference remains more efficient for larger models. To assess metric sensitivity, we recompute the correlation shown in Fig. 3. For this test, we use uncertainty weights derived from a single model, keeping weights fixed rather than adapting them per model. Specifically, we compute the overall correlation with the Driving Score using weights from a model trained for 10 epochs (achieving a 0.67 Pearson coefficient) and 40 epochs (Pearson coefficient of 0.73). These results indicate that UWE is not highly sensitive to model selection. While we evaluate UWE across models and deployment settings, including real-world applications on a physical platform, each metric introduces trade-offs. For instance, UWE does not require HD Maps or the meticulous offline simulation design used in NAVSIM [61].

### v2

@@ IV Experiments
We first describe our experimental setup and model training and testing procedure. We then leverage the setup to perform comprehensive experiments in simulation over various towns, ambient settings, and traffic infractions using the CARLA simulation [Dosovitskiy2017CORL]. However, while most online performance studies for autonomous driving are performed in simulation, this may bias results, i.e., due to appearance and physics artifacts. To ensure our findings generalize to real-world models, we thus also incorporate analysis using a small-scale platform in the real-world.
@@ IV-A Experimental Setup
Simulation Benchmark: We follow the standard Longest6 testing setup  [chitta2022transfuser, jaeger2023hidden]. Longest6 benchmark comprises 36 routes with an average route length of 1.5km, which goes beyond the prior study on CARLA’s Town02 by Codevilla [Codevilla2018ECCV]. We leverage a privileged rule-based expert to generate a large training and validation dataset, where trained policies are evaluated over unseen weather and routes in training.
@@
Real-World Benchmark: Online evaluation of driving policies is generally done in simulation [chitta2022transfuser, lbc, jiang2023vad]. To ensure our findings generalize to the real-world, we utilize a safe, small-scale platform. This platform is based on a commercial RC car (Traxxas X-Maxx [xmax]). We set up a Jetson Orin and a wide-angle camera to capture observations and control the steering and speed of the car using custom scripts and a joystick controller. We note that we do not leverage LiDAR in our real-world platform. We collect a training set of $32,356$ frames on various sidewalks, times of day, and weather. We evaluate on an uneven road that has three right turns, two left turns, and three pedestrians crossing (example routes are shown in Fig. 2). The platform can travel at speeds of up to 5050MPH. Online evaluations are safely done with an emergency software stop. In addition to the long route evaluation, we also perform targeted evaluations with clearly defined scenarios over traffic lights, vehicles, and object crossings.
@@
Model Training Strategies: In the simulation, our baseline model is TransFuser [jaeger2023hidden], which provides a state-of-the-art conditional imitation learning model. The model is a multi-modal policy with two inputs: RGB image and LiDAR. The features of these modalities are fused together by a transformer-based architecture with a cross-modal attention mechanism to predict output actions. To ensure our findings are applicable to a broad range of models, we train a large set of model variations by modifying the model input (e.g., RGB only vs. RGB and LiDAR) and backbone (two ConvNext [liu2022convnet], small and tiny, and nine RegNet [radosavovic2020designing], from RegNetY-200MF to RegNetY-8.0GF). We train each model (input and backbone combination) for 40 epochs while also saving the checkpoints every 10 epochs to sample additional model variations. Applying test-time dropouts provides another means of sampling models and covering the space of all possible policies. Moreover, we sample models using test-time dropout for values between 10% to 30% with increments of 10%. The Pearson correlation coefficient is then computed over all sampled models. Results are presented across 82 models, each varying in backbones, epochs, inputs, and test-time dropout rates. For real-world models, in the interest of optimizing for real-time performance, we target smaller and simpler architectures motivated by prior works from Codevilla et al. [Codevilla2018ICRA], Hawke et al. [hawke2020urban] and Bojarski et al. [bojarski2016end] with a baseline of a five-layer CNN as well as variants of RegNetY as feature extractor backbones. In the real-world experiment, we train an end-to-end policy model solely using RGB input, as it provides an easy and scalable implementation for researchers to reproduce and build upon in the future (i.e., without LiDAR).
@@
Figure 3: Driving Score Correlation Analysis in Simulation. The plot shows updated correlation for offline metrics, including TRE, QCE and PDM, the most successful reported metrics by prior research [Codevilla2018ECCV, dauner2024navsim], as well as widely employed metrics such as waypoint FDE (Final waypoint Displacement Error). Given our updated evaluation setup with complex traffic scenarios and diverse models, correlations are low overall, besides the introduced uncertainty-weighted (UW) version. Each disc shows one model sampled from a certain epoch, backbone, input (with or without LiDAR), and a test-time dropout rate used to obtain coverage of models with varying performances (see further explanation in Sec. IV-A). The radius of each marker is proportional to the percent of test-time dropout used (this sampling is independent of our proposed metric).
@@ IV-B Results in Simulation
Given the more extensive benchmark and model sampling strategies, i.e., with dropout over model weights for broader coverage of the performance spectrum, we revisit the analysis of Codevilla et al. [Codevilla2018ECCV] (only analyzed steering-based metrics, while fixing other action values based on the ground-truth). We also leverage the Driving Score, as it provides a holistic measure for driving quality, i.e., beyond Success Rate, as well as finer-grained online metrics.
@@
Revisiting Correlation in Simulation: As shown in Fig. 3, we show an overall poor correlation (both for Driving Score and Success Rate, isolated in Fig. 4) for standard offline metrics. Specifically, QCE and TRE, which are designed for higher correlation with the online Success Rate metric, show poor correlation (0.22 and 0.56, respectively) to more holistic online performance metrics such as the Driving Score. As shown in Fig. 3, QCE and TRE even perform worse than FDE (waypoint-based final displacement error), which is a standard offline metric for high-performing systems today. We hypothesize that the reason these were successful is due to the highly simplistic test settings in Codevilla et al. [Codevilla2018ECCV] (Town02, not considering collisions or infractions). However, we observe that our proposed UWE achieves the highest correlation within this challenging benchmark, $0.69$, $0.08$ higher than prior offline metrics. We only show selected metrics, as the remaining analyzed offline metrics performed with worse correlation. We study nine online metrics to get a complete understanding of the strengths and weaknesses of these metrics and identify key aspects that help choose the most promising metric. Our UWE demonstrates a stronger correlation with all online metrics and significantly outperforms other common offline metric baselines, as shown in Fig. 4. We also compare our UWE metric with the PDM (Predictive Driver Model) score from NAVSIM [dauner2024navsim], as shown in Fig. 3, using the publicly available code. As shown in Fig. 3, the PDM score is lower in our more realistic CARLA simulation compared to its performance on nuPlan. Specifically, in settings with more complex physics and reactivity, UWE’s correlation exhibits significantly higher fidelity than the PDM score. We note that this is likely because the PDM score makes assumptions about agent reactivity and environmental information (i.e., being overly accommodating). In contrast, UWE does not require such assumptions. As such, our metric provides a simple yet effective and scalable alternative potentially applicable to a broader range of situations with various agent dynamics.
@@ IV-D UWE Ablation
We further investigate the reliability of the uncertainty estimates obtained via MC dropout. When we replace the dropout mechanism with a model ensemble and observe an increase in correlation from 0.69 to 0.80 in the Pearson coefficient. While this aligns with prior literature [ovadia2019can], dropout-based inference remains more efficient for larger models. To assess metric sensitivity, we recompute the correlation shown in Fig. 3. For this test, we use uncertainty weights derived from a single model, keeping weights fixed rather than adapting them per model. Specifically, we compute the overall correlation with the DS using weights from a model trained for 10 epochs (achieving a 0.67 Pearson coefficient) and 40 epochs (Pearson coefficient of 0.73). These results indicate that UWE is not highly sensitive to model selection. While we evaluate UWE across models and deployment settings, including real-world applications on a physical platform, each metric introduces trade-offs. For instance, UWE does not require HD Maps or the meticulous offline simulation design used in NAVSIM [dauner2024navsim].